# Causality, Emergence, Causal Emergence, Causal Emergence Identification, Causal Representation Learning, World Model

### 28 Dec 2023 v1

@@ Abstract
Emergence and causality are two fundamental concepts for understanding complex systems. They are interconnected. On one hand, emergence refers to the phenomenon where macroscopic properties cannot be solely attributed to the cause of individual properties. On the other hand, causality can exhibit emergence, meaning that new causal laws may arise as we increase the level of abstraction. Causal emergence theory aims to bridge these two concepts and even employs measures of causality to quantify emergence. This paper provides a comprehensive review of recent advancements in quantitative theories and applications of causal emergence. Two key problems are addressed: quantifying causal emergence and identifying it in data. Addressing the latter requires the use of machine learning techniques, thus establishing a connection between causal emergence and artificial intelligence. We highlighted that the architectures used for identifying causal emergence are shared by causal representation learning, causal model abstraction, and world model-based reinforcement learning. Consequently, progress in any of these areas can benefit the others. Potential applications and future perspectives are also discussed in the final section of the review.
@@
Causality, Emergence, Causal Emergence, Causal Emergence Identification, Causal Representation Learning, World Model
@@ 1 Introduction
How did the first living cell emerge from the collisions between various molecules in the early Earth’s environment Artime and De Domenico (2022)? How does the cognitive concept of "I" emerge from the intricate interactions among countless neurons in our brain Lagercrantz and Changeux (2009)? How do large neural language models suddenly exhibit emergent abilities Wei et al. (2022)? These fundamental questions revolve around the concept of emergence in complex living, cognitive, and artificial systems. Emergence refers to the phenomenon where macroscopic properties and phenomena that cannot be solely attributed to or explained by the properties of individual components Anderson (1972); Meehl and Sellars (1956); Bedau (1997); Holland (2000). This presents a formidable challenge to the traditional reductionist perspective, while also shedding light on the underlying reasons for the enigmatic nature of emergent phenomena.
However, as elucidated by Bedau’s theory of weak emergence Bedau (1997), many emergent phenomena can be comprehended through the interactions among the individuals within the system Holland (2000). Complex systems, in fact, consist of extensive networks of interacting components Albert and Barabási (2002). Within these networks, even a minor cause, such as a perturbation of a single unit, can propagate through the interconnected network, resulting in a collective effect. The phenomenon described is commonly known as the butterfly effect Wikipedia (2023a), which provides an explanation for the occurrence of emergence. On the other hand, emergent properties, such as homeostasis Wikipedia (2023b), can stabilize the system itself, thereby preserving the locality of causal effects and preventing the observation of macro-level effects. These phenomena demonstrate that complex systems achieve interactions through causal laws, where numerous local causal laws form interconnected causal networks as a whole. And this whole possesses its unique causal characteristics.
@@
Emergence and causation interconnect with each other. On the one hand, emergence is the causal effect of complex and nonlinear interactions between components in complex systems Anderson (1972); Meehl and Sellars (1956). On the other hand, emergent properties may have causal effects on individuals in complex systems Kim (1992); Holland (2000). For example, the price of fossil fuel is the emerging result of the interactions between buyers and sellers in the market, while the price may also give feedback to the market: it may affect the decision making of each individual.
@@
Furthermore, we can understand emergence through the perspective of causation. What does emergence mean is that some phenomena and properties on the macroscopic level can not be attributed to the microscopic properties Fromm (2005). Thus, emergent properties or phenomena lose their direct explanations as usual, but may be attributed to the causes on the macro-level as pointed out by Hoel et al. (2013). Therefore, new causalities can be observed on larger scales.
@@
In conclusion, gaining a deep understanding of emergence is crucial in the field of complex system studies. Specifically, the development of a quantitative theory of emergence is on the verge of emerging. Such a theory holds the potential to address significant challenges, including the origins of life, the emergence of novel capabilities in large neural network models, and the potential for intelligence, consciousness, and free will to arise in artificial systems. Causality not only exhibits a profound connection with emergence but is also considered by many researchers as one of the most crucial perspectives for quantitatively comprehending emergence Hoel et al. (2013); Crutchfield and Young (1989); Fromm (2005); Seth (2008).
@@
Two primary challenges take precedence in understanding emergence from a causal perspective. The first is establishing a quantitative definition of emergence, while the second involves identifying emergent behaviors or phenomena through data analysis.
@@
To address the first challenge, two prominent quantitative theories of emergence have emerged in the past decade. The first is Erik Hoel’s theory of causal emergence Hoel et al. (2013), while the second is Rosas’ theory of emergence based on partial information decomposition Rosas et al. (2020).
@@
Hoel’s theory of causal emergence specifically addresses complex systems that are modeled using Markov chains. It employs the concept of effective information (EI) to quantify the extent of causal influence within Markov chains and enables comparisons of EI values across different scales Hoel et al. (2013); Hoel (2017). Causal emergence is defined by the difference of EI values between macro-level and micro-level. Several examples of discrete Markov chains have been shown to exhibit causal emergence if their EI values of macro-level dynamics are larger than the micro-level. Hoel and other researchers further extended measures of effective information and causal emergence for dynamical systems with continuous variables in Chvykov and Hoel (2020) and complex networks Klein and Hoel (2020). Other measures are also possible to quantify causal effects and thereafter causal emergence. In Comolatti and Hoel (2022), Hoel systematically compared a bunch of causal effect measures and concluded that causal emergence is independent on the selection of the measure.
@@
However, it is essential to establish a coarse-graining strategy beforehand in Hoel’s theory of causal emergence. Alternatively, the strategy can be derived by maximizing the effective information (EI) Hoel et al. (2013). However, this task becomes challenging for large-scale systems due to the computational complexity involved. To address these problems, Rosas introduced a new quantitative definition of causal emergence Rosas et al. (2020) that does not depend on coarse-graining methods, drawing from Partial Information Decomposition (PID) related theory. PID is an approach developed by Williams et al., which seeks to decompose the mutual information between a target and source variables into non-overlapping information atoms: unique, redundant, and synergistic information Williams and Beer (2010). Based on this groundwork, Rosas further developed the concept and introduced a theory called $\phi$ID to decompose the mutual information between multiple target and source variables Mediano et al. (2019). This framework provides a quantitative definition of causal emergence by measuring the positive synergy information between source and target variables based on the inherent characteristics of the system.
The second challenge pertains to the identification of emergence from data. In an effort to tackle this issue, Rosas et al. have derived a numerical method Rosas et al. (2020). However, it is important to acknowledge that this method offers only a sufficient condition for emergence and is an approximate approach. Another limitation is a coarse-grained macro-state variable should be given beforehand to apply this method. Hence, there is a need for the development of new methods. In a recent study by Zhang et al. Zhang and Liu (2022), a machine learning framework named Neural Information Squeezer (NIS) was introduced to tackle the challenge of identifying causal emergence using Hoel’s definition. Remarkably, the NIS neural network, functioning as a "machine observer" equipped with an internal model, exhibits a remarkable ability to identify causal emergence across various types of data. This aligns with the viewpoint suggested by Eric Bonabeau and Jean-Louis Dessalles Bonabeau and Dessalles (2011) that all emergent properties and phenomena are observer-dependent, and the epistemological nature of emergent causal powers, including downward causation Kim (1992); Crutchfield and Young (1989). However, employing a machine observer offers additional advantages in identifying emergence from data, as it helps avoid subjective judgments.
Furthermore, we find that causal emergence identification and the recent development of artificial intelligence can have deeper connections through the bridge of NIS and its extended frameworks Yang et al. (2023). First, the architecture of NIS, that is, the “encoder - dynamics learner - decoder” framework, is similar to causal representation learning Schölkopf et al. (2021) in artificial intelligence where dynamics learner is replaced as causal mechanisms. Therefore, there is a deep connection between the two emerging fields: causal emergence and causal representation learning. The former tries to characterise and understand the emergent causality in a complex system and the latter tries to learn the latent causal structures and mechanisms from observational data. They are connected by the problem of causal emergence identification. That is to identify causal emergence from data is essentially equivalent to learning the latent causality inherent in the data.
Second, there are more connections between causal emergence and other areas in artificial intelligence. For example, causal model abstraction also discussed the interplay between causality and the abstraction (coarse-graining) of causal models Beckers et al. (2020). In reinforcement learning based on world models Ha and Schmidhuber (2018), the similar framework, “encoder - dynamics learner - decoder”, is also used. Thus, the similarity and the shared features between the two fields can help us to lend the ideas and techniques in one field to the other field. For instance, EI may be used to measure the causal effect of neural networks Marrow et al. (2020), and the techniques in causal representation learning can be employed to detect and identify causal emergence. Zhang and Liu (2022).
@@
This paper tries to review the latest works on the quantitative theory and applications of causal emergence, and its relative works. We also discuss its connections with artificial intelligence and complex systems. In the second part, we introduce the background of causal emergence, particularly on the connections between causation and emergence in complex systems. The third part introduces several quantitative theoretic frameworks including: Crutchfield et al’s computational mechanics Crutchfield and Young (1989), Seth et al’s granger causal emergence Seth (2008), Hoel et al’s causal emergence theory and Rosas et al’ theory of emergence based on information decomposition. Several related conceptions are discussed, such as coarse-graining strategy, measure of effective information, and partial information decomposition. We also compare these two theories from several aspects. The forth part discusses the problem of causal emergence identification and then introduces the connection with artificial intelligence. Causal representation learning Schölkopf et al. (2021), abstraction of causal models Beckers et al. (2020), and world models in reinforcement learning Ha and Schmidhuber (2018) are introduced and the relations with causal emergence identification are discussed. Finally, we discuss the connections between causal emergence and other important topics in complex systems such as the origins of life, consciousness, etc.
@@ 2.1 Causality
Causality, which is a fundamental concept in many fields, including philosophy, natural science, and social science, refers to the relationship between a pair of events, where the first event (cause) can influence the second one (effect) Granger (1969); Pearl et al. (2000); Pearl (2009). The relationship between cause and effect can exhibit either deterministic or probabilistic characteristics. In deterministic causality, the cause will always produce the same effect, while in probabilistic causality, the cause will only produce the effect with a certain probability.
@@
In most physical systems governed by differential equations or Markovian dynamics, whether deterministic or probabilistic, causality is inherent. This arises from the fact that manipulating a variable within the system can result in observable changes in other variables, albeit with probabilistic outcomes. While, one physical mechanism may correspond several possible causal structures as pointed by Y. Iwasaki and H. A. Simon Iwasaki and Simon (1994) because the causal graphs describe local causal interactions while mechanisms are the global constraints for all variables.
Figure 1: The causal models (left column) and the corresponding causal hierarchy (right column) and .
@@
However, if we do not know the physical mechanism of a system, how can we model the causality behind the system? There are several parallel theoretical frameworks to do this. For example, Judea Pearl tries probabilistic graphic models (e.g. Bayesian network, causal graph, and structural causal model) to characterize causal interactions Pearl et al. (2000); Pearl (2009). In which nodes represent random variables, acyclic links represent causal interactions. Three different causal hierarchies are distinguished and quantified in different models. The first level of causality is association, i.e., the correlation between variables, which can be modelled by Bayesian networks Sucar (2015). The second level is intervention, which can be characterised by causal graph models Pearl (2009). The difference between a causal graph and a Bayesian network is the former allows defining a special operator which is called "do" operator to simulate the interventions done by the experimenter. The last level is courterfactual, which can be quantified by structural causal models Pearl et al. (2000). The distinguishing characteristic of a structural causal model is that all interactions within a causal graph can be described by a set of deterministic functions that involve the input variables and unknown components, thereby introducing uncertainty. Counterfactual causal inferences, that is the inferences under an imaginary scenario, can be conducted only on structural causal models. The different types of causal models and the corresponding causal hierarchy in Pearl’s theory are shown in Figure 1.
@@
Another quantitative framework for causal inference is Donald B Rubin’s "potential outcome" theory which is directly defined on the level of counterfactual Rubin (2005). According to this theory, each individual has two potential outcomes: the outcome that would be observed when individuals receive the treatment and the outcome that would be observed when they do not receive the treatment. Rubin’s theory suggests that the causal effect of a treatment can be estimated by comparing the difference between the observed outcome and the unobserved potential outcome for each individual.
@@
Researchers also developed a bundle of methods, which are named causal discovery algorithms Malinsky and Danks (2018), to automatically discover the causal relations between variables merely from data. For example, constraint-based algorithms Spirtes et al. (2000) and score-based algorithms Chickering (2002) for causal discovery are two typical representations . Causal discovery has become a hot interdisciplinary area in the middle of machine learning, causal inference, and statistics.
@@
Due to the uncertainty and ambiguity behind the discovered causal relations, measuring the degree of causal effect between two variables is another important problem. Numerous independent historical studies have tackled the issue of causation measurement. These include Hume’s concept of constant conjunction Hume (2016), Eells’ and Suppes’ measures of causation as probability raising Eells (1991); Suppes (1973), and Judea Pearl’s measures of causation Pearl (2009), etc. The effective information (EI) in the framework of causal emergence is also an indicator to measure causal effect. All these measures can almost be decomposed into two terms which characterize the necessary causation and sufficient causation, respectively Hoel and Levin (2020).
@@
In summary, causality is a fascinating inter-disciplinary topic across for several subjects. Quantitative framework and measures have been developed in last decades.

### 25 Feb 2024 v3

# causality; emergence; causal emergence; causal emergence identification; effective information; machine learning

@@ Abstract
Emergence and causality are two fundamental concepts for understanding complex systems. They are interconnected. On one hand, emergence refers to the phenomenon where macroscopic properties cannot be solely attributed to the cause of individual properties. On the other hand, causality can exhibit emergence, meaning that new causal laws may arise as we increase the level of abstraction. Causal emergence (CE) theory aims to bridge these two concepts and even employs measures of causality to quantify emergence. This paper provides a comprehensive review of recent advancements in quantitative theories and applications of CE. It focuses on two primary challenges: quantifying CE and identifying it from data. The latter task requires the integration of machine learning and neural network techniques, establishing a significant link between causal emergence and machine learning. We highlight two problem categories: CE with machine learning and CE for machine learning, both of which emphasize the crucial role of effective information (EI) as a measure of causal emergence. The final section of this review explores potential applications and provides insights into future perspectives.
@@
causality; emergence; causal emergence; causal emergence identification; effective information; machine learning
@@ 1 Introduction
How did the first living cell emerge from the collisions between various molecules in the Earth’s early environment Artime and De Domenico (2022)? How does the cognitive concept of “I” emerge from the intricate interactions among countless neurons in our brain Lagercrantz and Changeux (2009)? How do large neural language models suddenly exhibit emergent abilities Wei et al. (2022)? These fundamental questions revolve around the concept of emergence in complex living, cognitive, and artificial systems. Emergence refers to the phenomenon where macroscopic properties and phenomena cannot be solely attributed to or explained by the properties of individual components Anderson (1972); Meehl and Sellars (1956); Bedau (1997); Holland (2000). This presents a formidable challenge to the traditional reductionist perspective while also shedding light on the underlying reasons for the enigmatic nature of emergent phenomena.
However, as elucidated by Bedau’s theory of weak emergence Bedau (1997), many emergent phenomena can be comprehended through the interactions among the individuals within the system Holland (2000). Complex systems, in fact, consist of extensive networks of interacting components Albert and Barabási (2002). Within these networks, even a minor cause, such as a perturbation of a single unit, can propagate through the interconnected network, resulting in a collective effect. The phenomenon described is commonly known as the butterfly effect Wikipedia (2023a), which provides an explanation for the occurrence of emergence. On the other hand, emergent properties, such as homeostasis Wikipedia (2023b), can stabilize the system itself, thereby preserving the locality of causal effects and preventing the observation of macro-level effects. These phenomena demonstrate that complex systems achieve interactions through causal laws, where numerous local causal laws form interconnected causal networks as a whole. And this whole possesses unique causal characteristics.
@@
Emergence and causation interconnect with each other. On one hand, emergence is the causal effect of complex and nonlinear interactions between components in complex systems Anderson (1972); Meehl and Sellars (1956). On the other hand, emergent properties may have causal effects on individuals in complex systems Kim (1992); Holland (2000). For example, the price of fossil fuel is the emerging result of the interactions between buyers and sellers in the market. At the same time, the price may also provide feedback to the market: it can affect the decision making of each individual.
@@
Furthermore, we can understand emergence through the perspective of causation. What emergence means is that some phenomena and properties on the macroscopic level can not be attributed to the microscopic properties Fromm (2005). Thus, emergent properties or phenomena lose their direct explanations as usual but may be attributed to the causes on the macro-level, as pointed out by Hoel et al. (2013). Therefore, new causalities can be observed on larger scales.
@@
In conclusion, gaining a deep understanding of emergence is crucial in the field of complex system studies. Specifically, the development of a quantitative theory of emergence is on the verge of emerging. Such a theory holds the potential to address significant challenges, including the origins of life Ellis (2023), the emergence of novel capabilities in large neural network models Wei et al. (2022), and the potential for intelligence, consciousness, and free will to arise in artificial systems Oizumi et al. (2014). Causality not only exhibits a profound connection with emergence but is also considered by many researchers as one of the most crucial perspectives for quantitatively comprehending emergence Hoel et al. (2013); Crutchfield and Young (1989); Fromm (2005); Seth (2008).
@@
Two primary challenges take precedence in understanding emergence from a causal perspective. The first is establishing a quantitative definition of emergence, whereas the second involves identifying emergent behaviors or phenomena through data analysis.
@@
To address the first challenge, two prominent quantitative theories of emergence have emerged in the past decade. The first is Erik Hoel et al.’s theory of causal emergence Hoel et al. (2013), whereas the second is Fernando E. Rosas et al.’s theory of emergence based on partial information decomposition Rosas et al. (2020).
@@
Hoel et al.’s theory of causal emergence specifically addresses complex systems that are modeled using Markov chains. It employs the concept of effective information (EI) to quantify the extent of causal influence within Markov chains and enables comparisons of EI values across different scales Hoel et al. (2013); Hoel (2017). Causal emergence is defined by the difference in the EI values between the macro-level and micro-level. Several examples of discrete Markov chains have been shown to exhibit causal emergence if their EI values of macro-level dynamics are larger than those of micro-level dynamics. Hoel and other researchers further extended the measures of effective information and causal emergence for dynamical systems with continuous variables Chvykov and Hoel (2020) and complex networks Klein and Hoel (2020). Other measures are also possible for quantifying causal effects and, consequently, causal emergence. In Comolatti and Hoel (2022), Comolatti and Hoel systematically compared several causal effect measures and concluded that causal emergence is independent of the selection of the measure.
@@
However, in Hoel’s theory of causal emergence, it is essential to establish a coarse-graining strategy beforehand. Alternatively, the strategy can be derived by maximizing the effective information (EI) Hoel et al. (2013). However, this task becomes challenging for large-scale systems due to the computational complexity involved. To address these problems, Rosas et al. introduced a new quantitative definition of causal emergence Rosas et al. (2020) that does not depend on coarse-graining methods, drawing from partial information decomposition (PID)-related theory. PID is an approach developed by Williams et al., which seeks to decompose the mutual information between a target and source variables into non-overlapping information atoms: unique, redundant, and synergistic information Williams and Beer (2010). Based on this groundwork, Rosas further developed the concept and introduced a theory called $\phi$ID to decompose the mutual information between multiple target and source variables Mediano et al. (2019). This framework provides a quantitative definition of causal emergence by measuring the positive synergy information between the source and target variables based on the inherent characteristics of the system.
The second challenge pertains to the identification of emergence from data. In an effort to address this issue, Rosas et al. derived a numerical method Rosas et al. (2020). However, it is important to acknowledge that this method offers only a sufficient condition for emergence and is an approximate approach. Another limitation is that a coarse-grained macro-state variable should be given beforehand to apply this method. Hence, there is a need for the development of new methods.
Recently, artificial intelligence, propelled by the rapid advancements in machine learning and deep neural network technology, has witnessed significant progress. In the context of causal emergence, there are two key aspects to consider. Firstly, machine learning and neural network technology can be employed to address the challenge of identifying causal emergence. By leveraging these tools, we can develop approaches to effectively detect and analyze causal emergence phenomena. Secondly, the concepts and techniques from causal emergence can be introduced into machine learning to enhance the generalization capabilities of models. This integration can potentially improve the ability of machine learning algorithms to generalize well beyond the training data, leading to more robust and adaptable systems.
In a recent study by Zhang et al. Zhang and Liu (2022), a machine learning framework named the Neural Information Squeezer (NIS) was introduced to address the challenge of identifying causal emergence using Hoel et al.’s framework. Remarkably, the NIS neural network, functioning as a “machine observer” equipped with an internal model, exhibits a remarkable ability to identify causal emergence across various types of data. In the latest updated version of this work, the Neural Information Squeezer Plus (NIS+) has been developed to directly maximize the critical measure of causal emergence theory, namely effective information (EI) Yang et al. (2023). Through extensive experiments conducted on both simulated data and real brain data, the NIS+ has demonstrated its ability to automatically find emergent macro-variables and macro-dynamics. Consequently, the NIS+ enables quantifying causal emergence in data with the learned macro-dynamics. The results of these experiments highlight the effectiveness and potential of the NIS+ in capturing and analyzing causal emergence phenomena.
Furthermore, the NIS+ showcases superior performance in terms of generalization ability by directly maximizing effective information (EI). This brings forth a second question: can we leverage the measure of causation, EI, in the context of causal emergence, to enhance the generalization capability of neural networks for out-of-distribution scenarios? This concept is referred to as causal emergence for machine learning. By exploring this idea, we aim to bridge the gap between causal emergence and machine learning, potentially unlocking new avenues for improving the generalization abilities of machine learning.
Finally, in Section 5, we address several important and related issues. Firstly, we explore the similarities and differences between two emerging fields: causal emergence and causal representation learning Schölkopf et al. (2021). This comparison sheds light on the interplay between these two domains. Secondly, we delve into a philosophical problem concerning ontological or epistemological causality and emergence, providing insights into the underlying philosophical implications. Lastly, we discuss the potential applications of causal emergence in complex systems and how it contributes to our understanding of complex systems from a causal emergence perspective. These discussions broaden the scope of this paper and offer intriguing avenues for future research.
@@
This paper aims to provide a comprehensive review of the latest research on the quantitative theory and applications of causal emergence and related works. It also explores the connections between causal emergence, machine learning, and complex systems. The subsequent section delves into the background of causal emergence, with a particular focus on the interplay between causation and emergence in complex systems. In Section 3, various quantitative theoretical frameworks are introduced, including Crutchfield et al.’s computational mechanics Crutchfield and Young (1989), Seth et al.’s Granger causal emergence Seth (2008), Hoel et al.’s causal emergence theory, and Rosas et al.’s theory of emergence based on information decomposition. Additionally, related concepts such as the coarse-graining strategy, measures of effective information, and partial information decomposition are discussed, and a comparative analysis of these theories is presented. Section 4 addresses the connection between causal emergence theory and machine learning. It explores the use of machine learning and neural network techniques for identifying causal emergence and extends the measure of effective information (EI) to machine learning problems. Finally, this paper delves into other important topics and potential applications in the fields of machine learning and complex systems.
@@ 2.1 Causality
Causality, which is a fundamental concept in many fields, including philosophy, natural science, and social science, refers to the relationship between a pair of events, where the first event (cause) can influence the second one (effect) Granger (1969); Pearl et al. (2000); Pearl (2009). The relationship between cause and effect can exhibit either deterministic or probabilistic characteristics. In deterministic causality, the cause will always produce the same effect, whereas in probabilistic causality, the cause will only produce the effect with a certain probability.
@@
In most physical systems governed by differential equations or Markovian dynamics, whether deterministic or probabilistic, causality is inherent. This arises from the fact that manipulating a variable within the system can result in observable changes in other variables, albeit with probabilistic outcomes. As pointed out by Y. Iwasaki and H. A. Simon Iwasaki and Simon (1994), one physical mechanism may correspond to several possible causal structures because causal graphs describe local causal interactions, whereas mechanisms are the global constraints for all variables. We further discuss the relationship between Markovian dynamics and causal models in detail below.
@@
If we do not know the physical mechanism of a system, how can we model the causality behind the system? There are several parallel theoretical frameworks that can be used to do this. For example, Judea Pearl utilized probabilistic graphic models (e.g., Bayesian network, causal graph, and structural causal model) to characterize causal interactions Pearl et al. (2000); Pearl (2009). In these models, nodes represent random variables and acyclic links represent causal interactions. Pearl distinguished and quantified three levels of causality in different models. The first level is association, i.e., the correlation between variables, which can be modeled using Bayesian networks Sucar (2015). The second level is intervention, which can be characterized using causal graph models Pearl (2009). The difference between a causal graph and a Bayesian network is that the former enables defining a special operator, called a “do” operator, to simulate the interventions performed by the experimenter. The last level is counterfactual, which can be quantified using structural causal models Pearl et al. (2000). The distinguishing characteristic of a structural causal model is that all interactions within a causal graph can be described by a set of deterministic functions that involve the input variables and unknown components, thereby introducing uncertainty. Counterfactual causal inferences, that is, inferences under an imaginary scenario, can be conducted only on structural causal models. The different types of causal models and the corresponding causal hierarchy in Pearl’s theory are shown in Figure 1.
Figure 1: The causal models (left) column and the corresponding causal hierarchy (right) column.
@@
Another quantitative framework for causal inference is Donald B Rubin’s “potential outcome” theory, which operates at the counterfactual level Rubin (2005). According to this theory, each individual has two potential outcomes: the outcome that would be observed if the individual receives the treatment, and the outcome that would be observed if the individual does not receive the treatment. Rubin’s theory suggests that the causal effect of a treatment can be estimated by comparing the difference between the observed outcome and the unobserved potential outcome for each individual.
@@
Researchers have also developed a bundle of methods, named causal discovery algorithms Malinsky and Danks (2018), to automatically discover the causal relations between variables merely from data. For example, constraint-based algorithms Spirtes et al. (2000) and score-based algorithms Chickering (2002) for causal discovery are two typical representations. Causal discovery has emerged as an interdisciplinary area at the intersection of machine learning, causal inference, and statistics.
@@
Due to the uncertainty and ambiguity behind the discovered causal relations, measuring the degree of causal effect between two variables is another important problem. Numerous independent historical studies have addressed the issue of causation measurement. These include Hume’s concept of constant conjunction Hume (2016), Eells’ and Suppes’ measures of causation as probability raising Eells (1991); Suppes (1973), and Judea Pearl’s measures of causation Pearl (2009). Effective information (EI) in the framework for causal emergence is also an indicator for measuring causal effect, and we explain this point in Section 3.5.3. All these measures can be roughly decomposed into two causal primitives, which characterize the necessary causation and sufficient causation, respectively Hoel and Levin (2020).
It is crucial to differentiate between Markovian dynamics or dynamical mechanism models and the causal models introduced above. In current causal emergence theories Hoel et al. (2013); Rosas et al. (2020), Markovian dynamics, rather than causal models, including causal graphs and structural causal models, play a central role. There are several distinctions between Markovian dynamics and these causal models.
Firstly, in Markovian dynamics, the temporal evolution of one or multiple variables is typically described, with causal relationships between these variables extending across different points in time. Consequently, all variables are dependent on time, and a variable at one time step can causally impact another variable at the subsequent time step. Conversely, the aforementioned causal models do not explicitly incorporate the element of time. In these models, neither the variables nor the causal relationships are explicitly defined as functions of time.
Secondly, it is worth noting that circular interaction structures, including self-looped interactions, may exist in Markovian dynamics. However, such circular structures are typically not permitted in the causal models we introduced. This distinction arises from the fact that both causal graph models and structural causal models are constructed based on directed acyclic graphs (DAGs) Pearl (2009). However, this issue can be resolved by converting Markovian dynamics into DAGs. The initial step involves introducing static causal variables at each time step. Consequently, the same variable in Markovian dynamics can be transformed into a set of different variables at different time steps. This expansion allows for causal relationships across time steps while eliminating circular structures in causal graphs Koller and Friedman (2009).
Given the limitations of conventional causal models based on directed acyclic graphs (DAGs) Textor et al. (2016), there is a need for new representation methods. For instance, Richardson Richardson (2013) expanded the causal modeling approach to include DAGs. Spirtes Spirtes (2013) discussed DAG representations of feedback models and applied them to economic processes. Lacerda et al. also developed a method utilizing independent components analysis to discover cyclic causal models Lacerda et al. (2012). Furthermore, Forré and Mooij introduced a novel causal discovery method for nonlinear structural causal models with cycles and latent confounders Forré and Mooij (2018). George T.H. Ellison proposed a methodology that leverages domain knowledge to extract DAGs from temporal relationships among multiple variables Ellison (2023). Causal discovery methods can also be extended to temporal data by segregating various temporal data analysis tasks, such as classification, clustering, and prediction. Recently, Gong et al. provided a comprehensive review of these techniques in Gong et al. (2023). These advancements have the potential to broaden the range of available causal models.
@@
In summary, causality is a fascinating interdisciplinary topic spanning several subjects. Numerous quantitative frameworks and measures have been developed in recent decades.